An analyst for FoodMax estimates that the demand for its “Brand X” potato chips is given by: lnQXd = 12.14 – 2.8ln PX + 3.4PY + 0.7 ln AX where QX and PX are the respective quantity and price of a four-ounce bag of Brand X potato chips, PY is the price of a six-ounce bag sold by its only competitor, and AX is FoodMax’s level of advertising on brand X potato chips. Last year, FoodMax sold 7 million bags of Brand X chips and spent $0.42 million on advertising. Its plant lease is $2.1 million (this annual contract includes utilities) and its depreciation charge for capital equipment was $2.8 million; payments to employees (all of whom earn annual salaries) were $0.8 million. The only other costs associated with manufacturing and distributing Brand X chips are the costs of raw potatoes, peanut oil, and bags; last year FoodMax spent $2.8 million on these items, which were purchased in competitive input markets. Based on this information, what is the profit-maximizing price for a bag of Brand X potato chips? Instruction: Enter your response rounded to the nearest penny (two decimal places). $
Answer #1

Profit maximizing price is calculated as below by using formula-

P = (E/1+E) MC   here MC = 2.8/7 = 0.4

P = (-2.8/1+ (-2.8)) * 0.4 = $0.62

Hence, the profit-maximizing price for a bag of Brand X potato chips = $0.62